# TimeDART: A Diffusion Autoregressive Transformer for Self-Supervised Time Series Representation

### v3

@@ 4.2 Main Result
Forecasting
In the in-domain setting and after downstream fine-tuning for the forecasting task, TimeDART outperforms its competing baselines in most experimental scenarios. As shown in Table 2, TimeDART achieved the best results in $83.3\%$ of the 24 evaluation metrics, with the remaining results being second best. TimeDART further achieves a $6.8\%$ reduction in MSE compared to random initialization and $3\%$ compared to the state-of-the-art benchmarks, despite most methods showing noticeable gains over the random initialized model. Compared to supervised models like PatchTST and DLinear, TimeDART, even with random initialization, consistently performs competitively or better in most cases. After pre-training, it significantly surpasses these models, highlighting that the use of a vanilla Transformer as the backbone does not limit TimeDART’s performance. On the contrary, this choice underscores the robustness of the proposed architecture. The visualized results are provided in Appendix G to demonstrate our method’s effectiveness.
Classification
The experimental results for the classification task are shown in Table 4. To further explore the effectiveness of TimeDART’s representation learning, we also performed supervised fine-tuning on the classification task. The results in Table 4 show that after pre-training, TimeDART outperforms the baselines in terms of accuracy and F1 score in most experimental settings, even surpassing supervised methods specifically designed for classification tasks. Notably, after pre-training, TimeDART improves its average accuracy by approximately $5.7\%$ compared to random initialization. This suggests that through high-level autoregressive optimization and low-level noise-added patch reconstruction, TimeDART captures deep discriminative features of time series data, ensuring its effectiveness in downstream classification tasks.
TCN as Backbone
To evaluate the effectiveness of TimeDART with different backbones, we replaced the Transformer encoder with a causal TCN and removed the causal mask in downstream tasks for consistency with the forecasting task. As shown in Table 5, the causal TCN performs slightly worse than the Transformer encoder in forecasting tasks, likely due to the mismatch between the TCN and the Transformer denoising decoder used in pre-training. However, in the classification task, the TCN achieves competitive results and even outperforms the Transformer encoder with random initialization, suggesting the TCN’s stronger ability to capture discriminative features in time series data. Importantly, the causal TCN still outperforms its random initialization, demonstrating the adaptability of TimeDART’s self-supervised pre-training across different backbones.
Table 5: Performance of TCN as backbone. Average MSE and MAE from 4 different predicted windows for forecasting while accuracy and F1 score for classification task.
@@
Table 6: The results of ablation study. Average MSE and MAE from 4 different predicted windows for forecasting while accuracy and F1 score for classification task.
@@
Table 7: Hyperparameter sensitivity snalysis of noise steps and noise schedulers. Average MSE and MAE from 4 different predicted windows for forecasting while accuracy and F1 score for classification task. See Appendix F.1 for full results.
@@
1250 0.257 0.315 1250 0.135 0.247 1250 0.9258 0.9279
@@ 4.3 Ablation Study
We investigated the effectiveness of two key modules: autoregressive mechanism and the denoising diffusion process. Four experimental settings were considered: TimeDART (the original model), $w/o$ AR (autoregressive mechanism removed), $w/o$ diff (denoising diffusion process removed), and $w/o$ AR-diff (both modules removed). Specifically, in the autoregressive removal experiment, we eliminated both the causal mask in the Transformer encoder and the mask in the denoising decoder. The results in Table 6 demonstrate that both the autoregressive generation and the denoising diffusion model play crucial roles in the effectiveness of this approach. Notably, removing the autoregressive mechanism results in performance even worse than random initialization, demonstrating its crucial role in capturing long-term dynamic evolution. Similarly, removing the denoising diffusion process also leads to significant performance degradation, highlighting its effective contribution to capturing subtle local pattern information.
4.4 Hyperparameter Sensitivity Analysis
@@
We first investigate two key parameters through the forward process: the total number of diffusion steps $T\in\{750,1000,1250\}$ and the noise scheduler $\alpha(s)$, comparing cosine and linear schedules. The number of diffusion steps, with higher $T$ values increasing pre-training difficulty, and the noise scheduler, with the cosine scheduler offering smoother transitions than the linear one, both influence the recovery of clean patches. As shown in Table 7, the total number of noise steps does not significantly impact the difficulty of pre-training. However, the cosine noise scheduler performs substantially better than the linear scheduler while the latter sometimes even yields worse results than random initialization. This highlights the critical importance of the noise scheduler, as insufficiently smooth noise addition can result in significantly poorer outcomes.
@@
We then evaluate the impact of two parameters through the reverse process: the layer numbers of the denoising decoder and the patch length. The layer numbers, selected from $[0,1,2,3]$, reflects the relative size of the denoising decoder compared to the fixed 2-layer encoder, with 0 layers serving as the ablation case $w/o$ diff. As shown in Figure 5, allocating too many layers to the denoising decoder can result in under-training of the representation network due to an over-concentration of parameters in the denoising component. Additionally, we explored the impact of different mask strategies applied to the denoising decoder, as shown in Figure 4. We found that the self-only mask already contains all the necessary information for denoising, nd adding extra information actually hinders the denoising process. These results are detailed in Appendix F.2.
@@
On the other hand, the patch length controls the amount of local segment information each patch carries. Longer patch lengths are more effective for datasets like Electricity, which exhibit higher redundancy between consecutive data points, while shorter lengths are preferred for datasets with less redundancy, as they capture finer temporal dynamics. Figure 5 underscores the importance of adaptive patch length selection based on dataset characteristics.
@@ 5 Conclusion
In this paper, we introduced TimeDART, a novel pre-training framework for self-supervised time series representation learning that unifies two effective generative paradigms. By employing a causal Transformer encoder with a patch-based embedding strategy to model global trends and incorporating a denoising diffusion process to capture fine-grained local patterns, TimeDART effectively addresses both long-term dynamic evolution and subtle local features. Extensive experiments on benchmark datasets for time series forecasting and classification revealed that TimeDART consistently outperforms existing methods, confirming its effectiveness. With its unique approach, TimeDART has the potential to pave the way for future advancements in self-supervised representation learning for time series.
@@ Appendix A Dataset Descriptions
To evaluate the effectiveness of our proposed TimeDART, we conducted extensive experiments on 6 time series datasets for forecasting and 6 for classification. These datasets cover a variety of application scenarios, including power systems, transportation networks, and human activity. For detailed descriptions of the datasets and their respective divisions, please refer to Table 8.
@@
PEMS (4 subsets) (PeMS, )
@@ B.2 Implementation Details for Fair Experiment
All experiments were implemented using PyTorch (Paszke et al., 2017) and executed on a single NVIDIA RTX 4090 24GB GPU. To ensure the fairness of comparisons, we adopted a unified encoder as the representation network for all self-supervised methods, maintaining consistency in model-specific parameters. For supervised methods, we adhered to their original implementations without altering model-specific parameters, making adjustments only to non-core training parameters for alignment. For forecasting experiments, we applied the channel-independent configuration (Nie et al., 2022), ensuring consistent settings across methods.
@@
The representation network consists of 2 layers for most datasets, while for the Traffic dataset, it comprises 3 layers due to its higher complexity. The batch size was set to 16 for all datasets, except for Traffic and PEMS datasets, where it was reduced to 8 due to memory and time limitations. The ADAM optimizer (Kingma & Ba, 2017) was used with initial learning rates selected from $\{10^{-3},5\times 10^{-4},10^{-4}\}$, and L2 loss was employed for optimization. The pre-training process spanned 50 epochs. Sequence representation dimensions were chosen from $\{8,16,32,64,128\}$.
@@
For the forecasting task, a consistent look-back window length of $L=336$ and predicted window lengths $H\in\{96,192,336,720\}$ were applied to all datasets, except for the PEMS datasets, where $L=96$ and $H\in\{12,24,36,48\}$. Fine-tuning was performed for 10 epochs, with settings largely consistent with pre-training. The channel-independent configuration (Nie et al., 2022) was utilized, ensuring fairness in encoding time series data across all methods. Mean Squared Error (MSE) loss was employed as the loss function for forecasting tasks, ensuring consistency across experiments. To emphasize the contributions of pre-training, we also conducted experiments with Random Init., where the representation network was randomly initialized and directly fine-tuned without pre-training.
@@ C.3 Optimization Objective Derivation
$\mathcal{L}_{ideal}=\sum_{j=1}^{N}H(p_{\theta}(x_{j}^{0}),~{}q(x_{j}^{0}))=\sum_{j=1}^{N}\mathbb{E}_{q(x_{j}^{0})}[-\log p_{\theta}(x_{j}^{0})]$ (14)
However, since directly optimizing the exact marginal likelihood is intractable, we instead minimize the Evidence Lower Bound (ELBO), given by:
$\mathcal{L}_{ideal}\leq\mathcal{L}_{ELBO}=\sum_{j=1}^{N}\mathbb{E}_{q(x_{j}^{0:T})}\left[-\log\frac{q(x_{j}^{1:T}|x_{j}^{0})}{p_{\theta}(x_{j}^{0:T})}\right]$ (15)
Following a series of derivations (Luo, 2022), the final loss function is:
$\mathcal{L}_{diff}=\mathcal{L}_{ELBO}=\sum_{j=1}^{N}\mathbb{E}_{\epsilon,q(x_{j}^{0})}\left[||x_{j}^{0}-g(\hat{z}^{in}_{j},~{}f(\bm{z}^{in}_{1:j-1}))||^{2}\right],$ (16)

### v5

@@ 4.2 Main Result
Table 5: Performance of TCN as backbone. Average MSE and MAE from 4 different predicted windows for forecasting while Accuracy and Macro-F1 for classification task.
@@
Forecasting
In the in-domain setting and after downstream fine-tuning for the forecasting task, TimeDART outperforms its competing baselines in most experimental scenarios. As shown in Table 2, TimeDART achieves the best results in $83.3\%$ of the 24 evaluation metrics, with the remaining results being second best. TimeDART further achieves a $6.8\%$ reduction in MSE compared to random initialization and $3\%$ compared to the state-of-the-art benchmarks, despite most methods showing noticeable gains over the random initialized model. Compared to supervised models like PatchTST and DLinear, TimeDART, even with random initialization, consistently performs competitively or better in most cases. Moreover, it significantly surpasses these models after pre-training, highlighting that the use of a vanilla Transformer as the backbone does not limit TimeDART’s performance. In fact, this choice underscores the robustness of the proposed pre-training framework. The visualized results are provided in Appendix H to demonstrate the effectiveness of our method.
Table 6: The results of ablation study. Average MSE and MAE from 4 different predicted windows for forecasting while Accuracy and Macro-F1 for classification task.
@@
Classification
The experimental results for the classification task are shown in Table 4. To further explore the effectiveness of TimeDART’s representation learning, we also perform supervised fine-tuning on the classification task. The results in Table 4 show that after pre-training, TimeDART outperforms the baselines in terms of accuracy and F1 score in most experimental settings, even surpassing supervised methods specifically designed for classification tasks. Notably, after pre-training, TimeDART improves its average accuracy by approximately $5.7\%$ compared to random initialization. This suggests that through high-level autoregressive optimization and low-level noise-added patch reconstruction, TimeDART captures deep discriminative features of time series data, ensuring its effectiveness in downstream classification tasks.
Table 7: Hyperparameter sensitivity analysis of noise steps and noise schedulers. Average MSE and MAE from 4 different predicted windows for forecasting while Accuracy and Macro-F1 for classification task. See Appendix F.1 for full results.
@@
1250 0.257 0.315 Random Init. 0.269 0.323 1250 0.135 0.247 Random Init. 0.145 0.255 1250 0.9258 0.9279 Random Init. 0.9265 0.9237
TCN as Backbone
To evaluate the effectiveness of TimeDART with different backbones, we replace the Transformer encoder with a causal TCN and remove the causal mask in downstream tasks for consistency. As shown in Table 5, the causal TCN performs slightly worse than the Transformer encoder in forecasting tasks, likely due to the mismatch between the TCN and the Transformer denoising decoder used in pre-training. However, in the classification task, TCN performs competitively, even outperforming the randomly initialized Transformer encoder. This suggests TCN’s stronger ability to capture discriminative time-series features. Crucially, TCN still surpasses its random initialization, demonstrating TimeDART’s adaptable self-supervised pre-training across diverse backbones.
@@ 4.3 Ablation Study
We investigate the effectiveness of two key modules: autoregressive mechanism and the denoising diffusion process. Four experimental settings are considered: TimeDART (the original model), $w/o$ AR (autoregressive mechanism removed), $w/o$ diff (denoising diffusion process removed), and $w/o$ AR-diff (both modules removed). Specifically, in the autoregressive removal experiment, we eliminate both the causal mask in the Transformer encoder and the mask in the denoising decoder. The results in Table 6 demonstrate that both the autoregressive generation and the denoising diffusion model play crucial roles in the effectiveness of this approach. Notably, removing the autoregressive mechanism results in performance even worse than random initialization, demonstrating its crucial role in capturing long-term dynamic evolution. Similarly, removing the denoising diffusion process also leads to significant performance degradation, highlighting its effective contribution to capture subtle local pattern information.
4.4 Hyperparameter Sensitivity
@@
We first investigate two key parameters through the forward process: the total number of diffusion steps $T\in\{750,1000,1250\}$ and the noise scheduler $\alpha(s)$, comparing cosine and linear schedulers. The number of diffusion steps, with higher $T$ values increasing pre-training difficulty, and the noise scheduler, with the cosine scheduler offering smoother transitions than the linear one, both influence the recovery of clean patches. As shown in Table 7, the total number of noise steps does not significantly impact the difficulty of pre-training. However, the cosine noise scheduler performs substantially better than the linear scheduler while the latter sometimes even yields worse results than random initialization. This highlights the critical importance of the noise scheduler, as insufficiently smooth noise addition can result in significantly poorer outcomes.
@@
We then evaluate the impact of two parameters through the reverse process: the layer numbers of the denoising decoder and the patch length. The layer numbers, selected from $[0,1,2,3]$, reflects the relative size of the denoising decoder compared to the fixed $2$-layer encoder, with 00 layers serving as the ablation case $w/o$ diff. As shown in Figure 5, allocating too many layers to the denoising decoder can result in under-training of the representation network due to an over-concentration of parameters in the denoising component. Additionally, we explore the impact of different mask strategies applied to the denoising decoder, as shown in Figure 4. We find that the self-only mask already contains all the necessary information for denoising, and adding extra information actually hinders the denoising process. These results are detailed in Appendix F.2.
Table 8: Fine-tuning under few-shot and linear probing settings. Average MSE and MAE from 4 different predicted windows for forecasting while Accuracy and Macro-F1 for classification task. See Appendix G for full results.
ETTh2 PEMS04 HAR
(a) Few-shot (b) Linear Probing (a) Few-shot (b) Linear Probing (a) Few-shot (b) Linear Probing
Portion MSE MAE Method MSE MAE Portion MSE MAE Method MSE MAE Portion ACC. F1 Method ACC. F1
5% 0.365 0.392 TimeDART 0.354 0.391 5% 0.142 0.254 TimeDART 0.145 0.258 5% 0.8843 0.8901 TimeDART 0.8976 0.9005
10% 0.354 0.391 SimMTM 0.357 0.395 10% 0.140 0.250 SimMTM 0.148 0.260 10% 0.9014 0.9053 SimMTM 0.8732 0.8756
Full 0.346 0.387 TimeMAE 0.361 0.397 Full 0.133 0.245 TimeMAE 0.152 0.264 Full 0.9247 0.9286 TimeMAE 0.8858 0.8862
Ran. Full 0.358 0.396 Random Init. 0.368 0.401 Ran. Full 0.145 0.255 Random Init. 0.161 0.271 Ran. Full 0.8738 0.8723 Random Init. 0.8542 0.8578
@@
On the other hand, the patch length controls the amount of local segment information that each patch carries. Longer patch length is more effective for datasets like Electricity, which exhibit higher redundancy between consecutive data points, while shorter length is preferred for datasets with less redundancy, as they capture finer temporal dynamics. Figure 5 underscores the importance of adaptive patch length selection based on dataset characteristics.
4.5 Analysis Experiment
Few-shot Evaluation
In real-world applications, labeled time series data often starts with a limited number of observations. This can make it tough to train accurate models. To see how well TimeDART performs in these situations, we conduct few-shot evaluations using the ETTh2, PEMS04, and HAR datasets. Specifically, we train our models using only a small fraction (either 5% or 10%) of the available training data, a method consistent with previous research (Liu et al., 2024d). We then evaluate these models on the complete test set. The results of this evaluation are presented in Table 8 (a).
The results show that while fine-tuning with a small portion of samples yields weaker performance than fine-tuning with the full dataset on both forecasting and classification tasks, it still outperforms full-data supervised fine-tuning under random initialization in the vast majority of cases. This corroborates the strong capabilities of the pre-trained representation network. The results are detailed in Appendix G.
Linear Probing
As an important fine-tuning setting, we also experiment with the linear probing, where we fix the pre-trained encoder and only fine-tune the newly added projector at the end of the model. Table 8 (b) illustrates the performance compared to some of the compared baselines.
Our findings indicate that TimeDART’s performance in the linear probing setup exceed that of several baseline models, providing additional evidence for the efficacy of our pre-training framework. Concurrently, linear probing often matches or surpasses full-data supervised fine-tuning from random initialization (e.g., MSE 0.354 vs. 0.358), demonstrating the generalizability of TimeDART’s learned representations across various downstream tasks. The results are detailed in Appendix G.
Adapt to Extended-Length Input
The standard forecasting framework can often struggle when faced with extended input lengths, a challenge frequently attribute to the increased noise inherent in longer time series. However, TimeDART’s pre-training explicitly incorporates noise, effectively teaching the model to extract valuable representations even from noisy data. This intentional design allows TimeDART to achieve superior forecasting performance with longer look-back windows.
Furthermore, as the look-back window increases, we observe a consistent improvement in forecasting in Figure 6. This demonstrates that TimeDART successfully captures meaningful information from extended input length, and crucially, robustly mitigates the impact of increasing noise that often accompanies longer inputs.
Figure 6: Pre-training and Fine-tuning the model to the inputs with extended length. The MSE averaged from all predicted horizons is reported.
@@ 5 Conclusion
In this paper, we introduced TimeDART, a novel pre-training framework for self-supervised time series representation learning that unifies two effective generative paradigms. By employing a causal Transformer encoder with a patch-based embedding strategy to model global trends and incorporating a denoising diffusion process to capture fine-grained local patterns, TimeDART effectively addresses both long-term dynamic evolution and subtle local features. Extensive experiments on benchmark datasets for time series forecasting and classification reveal that TimeDART consistently outperforms compared methods, confirming its effectiveness. A series of extra analyses confirmed TimeDART’s suitability and robustness in numerous self-supervised settings. With its unique approach, TimeDART has the potential to pave the way for future advancements in self-supervised representation learning for time series.
Acknowledgments
This research was supported by grants from the National Natural Science Foundation of China (62337001), the grants of Provincial Natural Science Foundation of Anhui Province (No.2408085QF193), the Key Technologies R & D Program of Anhui Province (No. 202423k09020039) and the Fundamental Research Funds for the Central Universities (No. YD2150002501, No. WK2150110032).
@@ Appendix A Dataset Descriptions
To evaluate the effectiveness of our proposed TimeDART, we conducted extensive experiments on 6 time series datasets for forecasting and 3 for classification. These datasets cover a variety of application scenarios, including power systems, transportation networks, and human activity. For detailed descriptions of the datasets and their respective divisions, please refer to Table 9.
@@
PEMS (4 subsets) (PeMS, 2001)
@@ B.2 Implementation Details for Fair Experiment
All experiments were implemented using PyTorch (Paszke et al., 2017) and executed on a single NVIDIA RTX 4090 24GB GPU. To ensure the fairness of comparisons, we adopted a unified encoder as the representation network for all self-supervised methods, maintaining consistency in model-specific parameters. For supervised methods, we adhered to their original implementations without altering model-specific parameters, making adjustments only to non-core training parameters for alignment. For forecasting experiments, we applied the channel-independent configuration (Nie et al., 2023), ensuring consistent settings across methods.
@@
The representation network consists of 2 layers for most datasets, while for the Traffic dataset, it comprises 3 layers due to its higher complexity. The batch size was set to 16 for all datasets, except for Traffic and PEMS datasets, where it was reduced to 8 due to memory and time limitations. The ADAM optimizer (Kingma & Ba, 2014) was used with initial learning rates selected from $\{10^{-3},5\times 10^{-4},10^{-4}\}$, and L2 loss was employed for optimization. The pre-training process spanned 50 epochs. Sequence representation dimensions were chosen from $\{8,16,32,64,128\}$.
@@
For the forecasting task, a consistent look-back window length of $L=336$ and predicted window lengths $H\in\{96,192,336,720\}$ were applied to all datasets, except for the PEMS datasets, where $L=96$ and $H\in\{12,24,36,48\}$. Fine-tuning was performed for 10 epochs, with settings largely consistent with pre-training. The channel-independent configuration (Nie et al., 2023) was utilized, ensuring fairness in encoding time series data across all methods. Mean Squared Error (MSE) loss was employed as the loss function for forecasting tasks, ensuring consistency across experiments. To emphasize the contributions of pre-training, we also conducted experiments with Random Init., where the representation network was randomly initialized and directly fine-tuned without pre-training.
@@ C.3 Optimization Objective Derivation
$\mathcal{L}_{\text{ideal}}=\sum_{j=1}^{N}H(p_{\theta}(x_{j}^{0}),~{}q(x_{j}^{0}))=\sum_{j=1}^{N}\mathbb{E}_{q(x_{j}^{0})}[-\log p_{\theta}(x_{j}^{0})]$ (17)
Use the Evidence Lower Bound (ELBO) method and ignore the patch index subscript:
$L_{\text{ideal}}$ $\leq\mathbb{E}_{q(x^{0})}\left(\mathbb{E}_{q(x^{1:T}|x^{0})}\left[\log\frac{q(x^{1:T}|x^{0})}{p_{\theta}(x^{0:T})}\right]\right)$ (18)
$=\mathbb{E}_{q(x^{0:T})}\left[\log\frac{q(x^{1:T}|x^{0})}{p_{\theta}(x^{0:T})}\right]$ (19)
$:=L_{\text{diff}}$ (20)
Expand and use the Bayes formula:
$L$ $=\mathbb{E}_{q(x^{0:T})}\left[\log\frac{q(x^{1:T}|x^{0})}{p_{\theta}(x^{0:T})}\right]$ (21)
$=\mathbb{E}_{q(x^{0:T})}\left[\log\frac{\prod_{s=1}^{T}q(x^{s}|x^{s-1})}{p(x^{T})\prod_{s=1}^{T}p_{\theta}(x^{s-1}|x^{s})}\right]$ (22)
$=\mathbb{E}_{q(x^{0:T})}\left[\log\frac{q(x^{T}|x^{0})}{p(x^{T})}+\sum_{s=2}^{T}\log\frac{q(x^{s-1}|x^{s},x^{0})}{p_{\theta}(x^{s-1}|x^{s})}-\log p_{\theta}(x^{0}|x^{1})\right]$ (23)
The first term is a constant, the third term is the reconstruction loss, and only the second term is discussed:
$\mathbb{E}_{q(x^{0:T})}\sum_{s=2}^{T}\log\frac{q(x^{s-1}|x^{s},x^{0})}{p_{\theta}(x^{s-1}|x^{s})}$ $=\int\text{d}x^{0:T}~{}q(x^{0:T})\cdot\sum_{s=2}^{T}\log\frac{q(x^{s-1}|x^{s},x^{0})}{p_{\theta}(x^{s-1}|x^{s})}$ (25)
$=\sum_{s=2}^{T}\mathbb{E}_{q(x^{0},x^{s})}\left[D_{KL}(q(x^{s-1}|x^{s},x^{0})\|p_{\theta}(x^{s-1}|x^{s}))\right]$ (26)
Applying Bayes’ theorem again:
$q(x^{s-1}|x^{s},x^{0})=\mathcal{N}(x^{s-1},\tilde{\mu}^{s}(x^{s},x^{0}),\tilde{\beta}_{s}I)$ (27)
It can be assumed that the predicted distribution can also be expressed as:
$p_{\theta}(x^{s-1}|x^{s})=\mathcal{N}(x^{s-1};\mu_{\theta}(x^{s},s),\sigma_{s}^{2}I)$ (28)
According to the KL divergence of two Gaussian distributions with the same variance, we can get:
$L_{\text{diff}}=\mathbb{E}_{q(x^{0},x^{s})}\left[\frac{1}{2\sigma_{s}^{2}}\|\tilde{\mu}_{s}(x^{s},x^{0})-\mu_{\theta}(x^{s},s)\|\right]$ (29)
Since $\tilde{\mu}_{s}$ and $x^{0}$ are linearly related, we turn to predict the original space, so:
$L_{\text{diff}}=\mathbb{E}_{\epsilon,q(x^{0})}\left[||x^{0}-x^{out}||^{2}\right].$ (30)
Adding the cumulative sum of subscripts and the notation from the original paper, we get:
$\mathcal{L}_{\text{diff}}=\sum_{j=1}^{N}\mathbb{E}_{\epsilon,q(x_{j}^{0})}\left[||x_{j}^{0}-g(\hat{z}^{in}_{j},~{}f(z^{in}_{1:j-1}))||^{2}\right]$ (31)